You are considering a 10-year, $1,000 par value bond. Its coupon rate is 9%, and interest is paid semiannually. If you require an "effective" annual interest rate (not a nominal rate) of 6.09%, how much should you be willing to pay for the bond? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Given effective annual interest rate = 6.09%

So, APR compounded semiannually = n*(((1+EAR)^(1/n))-1) = 2*((1.0609^0.5) - 1) = 6%

Now using Financial calculator to calculate value of this bond,

Coupon is 9% of 1000 semiannually,

So annual coupon = $45

total period = 2*10 = 20

So putting Values in Financial calculator

PMT = 45

N = 20

I/Y = 6/2 = 3

FV = 1000

compute for PV, we get PV = $1223.16

So, current price of bond = $ 1223.16